OpenStudy (anonymous):

In the combustion of hydrogen gas, hydrogen reacts with oxygen from the air to form water vapor. Hydrogen+Oxygen->Water If you burn 39.7 g of hydrogen and produce 355g of water, how much oxygen reacted?

OpenStudy (anonymous):

I started my problem by writing a balanced equation of the reaction but now I haven't the slightest idea of what to do with it. 2H2+O2->2H20

OpenStudy (anonymous):

How many moles are in 39.7 g of Hydrogen? How many moles of hydrogen are in 355 g of water? How many moles of water are in 355 g of water? We need these three values to determine if the combustion process is stoichiometric.

OpenStudy (anonymous):

How would I figure the moles H/355g H20?

OpenStudy (anonymous):

isn't simply a matter of 355 - 39.7 g oxygen

OpenStudy (anonymous):

355 g of water = 355/18 moles

OpenStudy (anonymous):

Did you figure out the number of moles of hydrogen and water?

OpenStudy (anonymous):

Hydrogen+Oxygen->Water If you burn 39.7 g of hydrogen and produce 355g of water, how much oxygen reacted? so you have 355 grams of water in the end or the process Find the g/mol of H2O 355g H2O *(Oxygen g/mol/H2O g/mol) = grams of oxygen you reacted

OpenStudy (anonymous):

To find the g/mol of H2O simply (hydrogen g/mol)*2 + (g/mol of oxygen) = H2O g/mol

OpenStudy (anonymous):

355g H2O *(15.9994 g/mol/18.0153 g/mol) = 315.28 grams

OpenStudy (anonymous):

of Oxygen were reacted theoretically of course assuming all the oxygen was pushed forward to water.

OpenStudy (anonymous):

if you look at the formula you will notice the g/mol cross out and the H2O and the H2O cross out so all you are left with is Oxygen and grams

OpenStudy (anonymous):

Hope this helps

OpenStudy (anonymous):

355 - 39.7 = 315.3 g

OpenStudy (anonymous):

jimmy rep your answer is just as good as mine but what i have provided her is much more universal for dealing with problems like this that are a little more complex but yeah.

OpenStudy (anonymous):

It's helping but I'm getting confused. I know that there's 39.7mol H per 39.7g H You said to do 39.7g*2=79.4 79.4+(g/mol of oxygen) = H2O g/mol Now I'm just guessing 79.4+16=95.4 Sorry guys it's been a few years since I've done chemistry

OpenStudy (anonymous):

yea - i agree

OpenStudy (anonymous):

from equation and using approximate moles of reactants 2 H2 = 4 (4*1) 02 = 32 (2*16) 2 H20 = 36 (4+32) ratio 1:8:9 39.7 g of hydrogen reacts with 8*39.7 g oxygen giving 9* 39.7 g water

OpenStudy (anonymous):

Okay, now I get it. At first I solved it the same as Jimmy because of the Law of Conservation of Mass but I figured I should learn the proper way. Thank you to all who helped!!

OpenStudy (anonymous):

yw